Ship reports banner with links Home IODP-USIO daily/weekly reports

< August | Next expedition >

IODP Expedition 327

Juan de Fuca Ridge-Flank Hydrogeology

Daily Science Report for 4 September 2010

LOCATION: In transit to Victoria, British Columbia

SCIENCE UPDATE: The positioning beacon was recovered at 0118 hr. The drill string and bottom-hole assembly were on deck by 0530 hr. The rig was secured for transit and the thrusters were raised at 0930 hr. The JOIDES Resolution was underway at full speed at 1015 hr. The pilot boarded at 0752 hr and the first line ashore at Ogden Point, Victoria, was at 0836 hr.
Core description and lab measurements continued on the Site U1363 cores. Finalized the Expedition Summary and Site U1363 reports. Cleaned work areas. Weather on site: wind 24 knots gusting to 28 knots, seas 6 ft, swells 8 ft, air temperature 14°C, max. heave 1 m, max. pitch 2°, max. roll 3°.


Daily Science Report for 3 September 2010

LOCATION: Hole U1363D (Proposed Site GRB-3A; 47° 17.5759'N, 128° 1.7641'W; water depth 2689 m)
Hole U1363E (Proposed Site GRB-1A; 47° 17.427'N, 128° 1.995'W; water depth 2688 m)
Hole U1363F (Proposed Site GRB-1A; 47° 17.427'N, 128° 1.995'W; water depth 2688 m)
Hole U1363G (Proposed Site GRB-1A; 47° 17.427'N, 128° 1.995'W; water depth 2688 m)

SCIENCE UPDATE: Core U1363D-5X advanced to 2920 m (231 m below seafloor) and the drill string was pulled out of the seafloor at 0700 hr, ending Hole U1363D. Hole U1363E was spudded at 0840 hr and was drilled without coring to establish the depth to basement. The sediment/basement interface was confirmed at 36 mbsf at 1000 hr. The drill string was pulled out of the seafloor at 1125 hr, ending Hole U1363E. Hole U1363F was spudded at 1200 hr and Cores 1H to 4H advanced to 35 mbsf by 1600 hr. APCT3 temperature measurements were taken with Cores 3H and 4H. The drill string was pulled out of the seafloor at 1735 hr, ending Hole U1363F. The ship was offset and basement contact was again established at 17 mbsf by washing down. Hole U1363G was spudded at 2000 hr and Cores 1H to 3H advanced to 24.9 mbsf by 2250 hr. The true advance is closer to 17 mbsf as the last core recovered mostly flow-in material. An APCT3 temperature measurement was taken with Core 2H. The drill string was pulled out of the hole and was at 2634 m by midnight.
Core description, shipboard sampling, and lab measurements continued on the Site U1363 cores. Writing of Expedition Summary and Site U1363 reports in progress. Weather on site: wind 19 knots gusting to 21 knots, seas 5 ft, swells 6 ft, air temperature 16°C, max. heave 1 m, max. pitch 0.5°, max. roll 0.5°.


Daily Science Report for 2 September 2010

LOCATION: Hole U1363D (Proposed Site GRB-3A; 47° 17.5759'N, 128° 1.7641'W; water depth 2689 m)

SCIENCE UPDATE: The SET tool was removed from the outer core barrel at 0145 hr. It was decided to drill another hole at this location to recover the sediment/basement interface. The drill string was run to the seafloor and Hole U1363D was spudded at 0745 hr. Drilling without coring advanced to 2887 m (198 m below seafloor) by 1530 hr. Cores 2X-4X advanced to 221.6 mbsf by 2200 hr.
Core description, shipboard sampling, and lab measurements continued on the Site U1363 cores. Writing of Expedition Summary and Site U1363 reports in progress. Weather on site: wind 10 knots gusting to 12 knots, seas calm ft, swells 5 ft, air temperature 17°C, max. heave 1 m, max. pitch 1°, max. roll 0.75°.


Daily Science Report for 1 September 2010

LOCATION: Hole U1363C (Proposed Site GRB-3A; 47° 17.5759'N, 128° 1.7641'W; water depth 2689 m)
Hole U1363D (Proposed Site GRB-3A; 47° 17.5759'N, 128° 1.7641'W; water depth 2689 m)

SCIENCE UPDATE: Drilling without coring in Hole U1363C advanced to 2859 m (170 m below seafloor) by 0430 hr. Cores 2X-5X were cored from 170 to 202.4 mbsf, with SET temperature measurements taken before Core 2X, before Core 4X, and after Core 5X. At 1445 hr, the SET tool became stuck in the outer core barrel and attempts to work it free failed. The drill string had to be pulled out of the hole so the SET tool could be recovered. Operations at Hole U1363C ended at 1730 hr when the drill string cleared the seafloor. A second, failed attempt was made to pull the SET tool free between 1800 and 1930 hr. The drill string continued to be pulled up from the seafloor and the bottom-hole assembly was set at 40 m.
Core description, shipboard sampling, and lab measurements continued on the Site U1363 cores. Revision of Methods in progress. Review of Site U1362 reports continued. Weather on site: wind 4 knots gusting to 6 knots, seas calm ft, swells 6 ft, air temperature 16°C, max. heave 1 m, max. pitch 1.5°, max. roll 0.75°.




About the IODP JRSO | Expeditions | Data & core samples | Tools & laboratories | Publications | Education |
Newsroom | Meetings & port calls | Employment | Contact us | Search | Site map | People | Intranet | Home

Accesibility Statement | Plug-in download | Privacy statement | ODP

For comments or questions: webmaster@iodp.tamu.edu

Copyright IODP JRSO

Modified on Tuesday, 08-Nov-2011 08:45:21 CST.